# Multi-bandpass Photometry for Exoplanet Atmosphere Reconnaissance (MPEAR) with the Habitable Worlds Observatory (HWO) I. Differentiating Earth from Neptunes During Discovery

### v1

@@ Abstract
We describe a new algorithm designed to determine the best combination of broadband photometric observations for extracting maximum information from the first visit. It identifies degeneracies in the orbital configurations, fluxes, and noise, and determines optimal secondary photometry bands to reduce these. We demonstrate its application by comparing an Earth seen at quadrature with a cold and a warm Neptune at inclined orbits and varying phases, with comparable flux in the discovery bandpass centered at 500 nm (20% bandwidth).
@@ 1 Introduction
The characterization of habitable terrestrial exoplanets is one of the main goals of the major space agencies, as highlighted in the US Astro2020 Decadal Survey (NAP26141). Astro2020 recommended the pursuit of a technical and scientific study for the Habitable Worlds Observatory (HWO), an IR/O/UV (i.e., infrared/optical/ultraviolet) “high-contrast direct imaging mission with a target off-axis inscribed diameter of approximately 6 meters” which resembles but does not coincide with the previous HabEx (Habitable Exoplanet Observatory, Gaudi2020) and LUVOIR (Large UV/Optical/IR Surveyor, LUVOIR) concepts. In recent months, efforts have ramped up to perform architecture trade studies to eventually define an HWO concept, and the community has been heavily involved in proposing relevant science cases for the new Observatory. This is a critical task at this point in time since, as stated in the Decadal, “implementation should start in the latter part of the decade [2022-2032], with a target launch in the first half of the 2040s”, pending an independent review on the scientific and technological maturity. As the architecture for HWO is being developed, it is crucial to optimize the observing strategies for efficient target identification and characterization so that the mission requirements can be matched to the planned observations.
@@
A main exoplanet-related requirement for an HWO-like mission is to observe 30 - 50 terrestrial planets around nearby FGK stars over the span of 5 mission years, though the potential of a serviceable observatory is also discussed and would extend and enhance the nominal mission. Even with state-of-the-art coronagraphs, the spectroscopic characterization of planets whose flux is 10 billion times smaller than their host star (approximately the contrast we would need to detect an Earth-like planet around the Sun) would require days of observation (see e.g. 2025ApJ...990...48U). Furthermore, the presence of a coronagraph limits us to only observe a portion of the wavelength range at a time, thus increasing the required observing time for a full characterization of the planetary spectrum from the UV (i.e., ultraviolet) to the NIR (i.e. near-infrared).
@@
It is therefore essential to define optimal observing strategies for efficiently identifying which planetary discoveries are most likely to be habitable, and therefore of highest priority for spectroscopic follow-up. For the best targets, HWO will acquire spectroscopic data on the full wavelength range, and the current strategy baselined for target prioritization is through spectroscopy “decision tree” observation sequences (Young2024) which prioritize key molecules to detect that can distinguish different planet types. Bayesian retrievals have been used to identify what is the minimum and optimal wavelength ranges to perform observations for a variety of planetary atmospheres (see e.g., 2018AJ....155..200F; 2022AJ....163..299D; 2023AJ....166..157D; 2023AJ....166..129L; 2025arXiv250714771K; 2025ApJ...990...48U); these observational requirements have recently been used when simulating exo-Earth survey yields (2024AAS...24324906M; 2024SPIE13092E..5MM; Stark2024; Stark2).
@@
The potential of colors and broadband photometry to qualitatively characterize exoplanets is not new. 2003ASPC..294..595T argued that colors and low-resolution spectra in the VIS range would allow us to characterize atmospheric and surface spectral signatures, as well as determine physical parameters. On the other hand, Cahoy2010 performed a study on the albedo of gas giants and highlighted that phase-dependence and the presence of clouds could make this less straightforward and multiple observations would be needed to determine cloud distributions and orbital parameters. JKT2016 empirically calculated the optimal photometric bins to differentiate an Earth from a variety of other planetary classes. To achieve this, they identified the filters that maximized the distance in color-color space between the Earth and every other modeled planet. 2018AJ....156..158B explored the classification of gas giants through color information obtained with potential filter options for the Nancy Grace Roman Space Telescope, and the relation with metallicity, temperature, clouds, and phases. The authors performed a supervised classification multivariate analysis to assess if photometry would be sufficient to retrieve the underlying physics of an exoplanet. The analysis was done every combination of two up to five filters and found that at least three filters are needed to provide any classification. Smith2020 calculated a grid of potential rocky atmospheres and evaluated the potential for multiband photometry to effectively distinguish different classes of planetary atmospheres. They found the technique effective for atmospheres exhibiting strong absorption lines, but more challenging for Earth-like abundances and requiring large apertures to improve results.
Beyond the challenges of degeneracies in planetary atmospheres, distinguishing an Earth from other planetary types is made even more difficult when considering orbital geometry effects and unknown phase. As shown in 2018haex.bookE..98R, the reflected light spectra of different planets span orders of magnitude when positioned at the same semi-major axis and phase. However, degeneracies can appear when altering these two parameters, and atmospheric features might become less distinguishable. Specifically, 2018haex.bookE..98R highlighted a scenario where a Neptune-like planet and an Earth would hardly be differentiable with each other at a given phase, especially at lower resolutions and within the average expected noise for current instrumentation. Given the limited observing time available for future direct imaging missions, developing strategies to efficiently distinguish between these planetary types becomes crucial for prioritizing truly habitable worlds over larger, uninhabitable planets that happen to fall within similar brightness ranges.
@@
We provide methodology and results for every step in the algorithm: Section 3 describes the orbital modeling; Section 4 deals with the radiative transfer flux calculations; Section 5 provides details on the Exposure Time Calculator for HWO and the noise simulations; Section 6 explores the best filter combinations for a successful discrimination between an Earth and Neptune-like planets. We discuss the implications and limitations of this study in Section 7, and we provide a summary in Section 8.
2 Methodology Overview
@@
We test this algorithm considering the Earth on a face-on orbit (quadrature phase) compared to two template Neptune scenarios (which we call “Warm” and “Cold”). We will expand on different phases and more realistic scenarios in an upcoming study (see Section 7.4).
3 Orbital Calculations
@@ 3.1 Results
In Figure 2 we show how orbits that satisfy condition 1) are distributed in semi-major axis space. We highlight in green the models that are within the commonly assumed boundaries for the habitable zone (HZ; Kopparapu2013-HABITABLEZONESMAINSEQUENCE), and in red are the ones outside the HZ (note that planets interior to the HZ would have smaller angular separations than our assumed 0.1 arcsec and thus would not satisfy the orbital condition). We notice that about 65% of the orbits that fit the assumed target projected separation are within the HZ. A non-negligible 35% of the compatible orbits lie well beyond the HZ, with most within 5 AU. If we assume no prior knowledge on the observed system, we would have no way of knowing from the first orbit which of these configurations is the correct one. While we could assume that it is more likely that a planet at the prescribed 0.1 arcsec separation is explained by an orbit within the HZ, scenarios including planets outside the HZ cannot be excluded.
It must be noted that these results are highly dependent on the uncertainty of the projected separation. For this study, we made an arbitrary choice to ensure that the potential Earth would be located within the HZ of its star (which explains the higher percentage of planets in this region), but in reality these requirements might be more relaxed, especially for more distant targets. We also assumed that a planet could be potentially found at any inclination or semi-major axis, disregarding any information that could come from planet population occurrence rates or precursor observations. We address the limitations of these assumptions in Section 7.4.
4 Flux Calculation
@@
Atmosphere Reference kofman24 2025arXiv250813368P 2019ApJ...887..166H
Figure 3: Temperature and abundance profiles for template Earth and Neptunes atmospheres. Left: Earth atmosphere from kofman24 Center: Cold Neptune atmosphere from 2025arXiv250813368P Right: Warm Neptune atmosphere from 2019ApJ...887..166H.
To calculate planetary reflectance spectra, we use the Planetary Spectrum Generator (PSG) [^1] https://psg.gsfc.nasa.gov (Villanueva2018). PSG is a radiative transfer model suite hosting tools that are widely used to model not only confirmed exoplanets, but also Solar System planets and cometary objects. PSG can compute 1D and 3D radiative transfer models from any configuration and allows the flexibility to choose instrumental parameters for nearly any observatory. It can ingest billions of spectral lines of almost 1000 species across its spectral repositories (HITRAN, JPL, CDMS, GSFC-Fluor, ExoMol), allowing the user to produce detailed high resolution spectra.
@@
We therefore enabled the “disk sub-sampling” mode in PSG to produce planetary reflectances at different phase angles. In this mode, PSG divides the disk of the planet into multiple concentric regions of common incidence and emission angles, each with a specific weight based on the disk percentage the region occupies. For every bin, the radiative transfer equation is solved and the output flux is a weighted mean of the fluxes of each region. In our study, we assumed a sampling of $N=5$, which creates five bins for the incidence angle and five more for the emission angle. As these two angles change with phase, the total amount of regions that are created (as the possible combinations of the incidence and emission angle) varies, up to around 25 ($\approx N^{2}$). 2021AJ....162...30S discusses PSG’s disk sub-sampling algorithm in more depth.
@@
The algorithm we developed for this study is interfaced with an instance of PSG installed locally via a Docker container [^2] https://docs.docker.com/get-started/docker-concepts/the-basics/what-is-a-container/, enabling repeated and rapid calls to PSG (see Villanueva2018; 2022fpsg.book.....V, for more information on how to implement a “dockerized” PSG version). For every model to be calculated, the algorithm produces a configuration file that describes the desired input (including the atmospheric profile, the desired opacities to be used, the surface and aerosol scattering parameters, the stellar class and the orbital parameters) as well as the format of the desired output, and then reads in the results of the calculation.
@@
For the reference scenario (Earth at quadrature), we calculate the flux of an Earth twin orbiting a Sun-twin star at 1 AU. For the atmospheric profile for a standard Earth-like planet, we use an atmosphere based on kofman24, modeled using the Modern-Era Retrospective analysis for Research and Applications (MERRA-2) 3D climate database [^3] https://gmao.gsfc.nasa.gov/gmao-products/merra-2/. The 3D distribution of temperature, pressure, clouds and relevant molecules from MERRA-2 were ingested into the Global Emission Spectrum (GlobES, 2025A&C....5300982F) module of the Planet Spectrum Generator (PSG Villanueva2018)). In the MERRA-2 reanalysis database, the 3D abundance and presence of clouds was parameterized in a 2D cloud fraction field. PSG evaluates the spectra from the MERRA-2 input by running two spectral calculations, one considering a clear sky and one assuming a fully clouded atmosphere according to the ingested cloud particle abundances. The two spectra are combined following the cloud fraction.
@@
Surface coverage was adopted from the Moderate Resolution Imaging Spectroradiometer (MODIS) [^4] https://modis.gsfc.nasa.gov using five different surface reflectances from the United States Geological Survey (USGS) [^5] https://www.usgs.gov/labs/spectroscopy-lab/usgs-spectral-library. The data from this source were validated with narrow band photometric observations from DSCOVR (between 0.318-0.780 microns), as well as several other observations of Earth (2025arXiv250813368P). We show in Figure 3 and Table 1 the atmospheric and temperature profiles for the Earth, as well as some relevant input parameters.
@@
We then calculate the flux of all compatible orbit configurations from Section 3, assuming two different Neptune scenarios: if the semi-major axis is less than 3.5 AU (1167 out of 1435 Neptunes), then we assume a warm Neptune atmosphere; otherwise (268 out of 1435 Neptunes), we assume a cold Neptune atmosphere. The choice of the 3.5 AU threshold is motivated by estimates for the snow line (Martin2012, see e.g.,), which is a sensible physical separation between classes of Neptune-like planets with different atmospheric compositions and thermal structures. Beyond this boundary, a water or ammonia cloud deck is expected above $\sim$1 bar (2019ApJ...887..166H; Cahoy2010).
@@
We use fixed atmospheric templates for each Neptune type to isolate the effects of orbital geometry (phase angle and distance) on observed fluxes, allowing us to focus on how orbital parameters alone can create spectral confusion. We discuss the implications of this assumption in Section 7.4 and we will consider self-consistent chemistry in a future work.
@@
As warm Neptune template, we use one of the models calculated in 2019ApJ...887..166H. We selected a model for an H2/He-dominated warm Neptune at a semi-major axis 1.7 AU (a good medium between the edges of the warm Neptune subgrid), internal heat flux $T_{int}=60\ K$ to mimic a Neptune-like planet, gravity of $10\ m/s^{2}$ (close to that of Neptune), and 10x solar metallicity (similar to the metallicity found in observed warm Neptune-sized planets). The models were computed using ExoREL, which calculates pressure-temperature profiles using a grey-atmosphere approximation and taking into account the condensation of water and ammonia clouds in a self-consistent way.
The cold Neptune template was derived from all available measurements of the Solar System’s Neptune (2025arXiv250813368P). The cold Neptune continuum is based on data from the Hubble Space Telescope (Courtin1999), the European Southern Observatory (Karkoschka1998) and the NASA Infrared Telescope Facility (Irwin2022). The collection of observations provide the composite reflectance spectrum that was used as a reference to develop the best-fit reflectance model in PSG by adjusting atmospheric absorbers, aerosols, and cloud features to match the reference spectrum; additional details are available in 2025arXiv250813368P.
We simulate the “discovery” flux measurement by calculating the average photometric flux of the Earth at 500 nm assuming 20% bandwidth $F_{500,20\%}$. This translates into averaging the flux in the interval $500\cdot(1-10\%),500\cdot(1+10\%)]$ nm =[450,550]=[450,550] nm. Since the currently-assumed signal-to-noise ratio ($S/N$) for detection is 7 (Nemati2020-MethodDerivingOptical) at the detection wavelength range, the errorbar associated with the measurement $\Delta F_{500,20\%}=F_{500,20\%}/7$. The flux of a Neptune is then considered compatible if the averaged Neptune flux over the discovery bandpass lies within $F_{500,20\%}\pm\Delta F_{500,20\%}$.
4.1 Results

### v2

@@ Abstract
We describe a new open-source algorithm designed to determine the best combination of broadband photometric observations for extracting maximum information from the first visit. It identifies degeneracies in the orbital configurations, fluxes, and noise, and determines optimal secondary photometry bands to reduce these. We demonstrate its application by comparing an Earth seen at quadrature with a cold and a warm Neptune at inclined orbits and varying phases, with comparable flux in the discovery bandpass centered at 500 nm (20% bandwidth).
@@
software: MPEAR (Alei, 2026), pyEDITH (Alei et al., 2025b), spectres (Carnall, 2017), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), numpy (Harris et al., 2020), pandas (The pandas development team, 2020; McKinney, 2010), matplotlib (Hunter, 2007).
@@ I Introduction
The characterization of habitable terrestrial exoplanets is one of the main goals of the major space agencies, as highlighted in the US Astro2020 Decadal Survey (National Academies of Sciences, Engineering, and Medicine, 2021). Astro2020 recommended the pursuit of a technical and scientific study for the Habitable Worlds Observatory (HWO), an IR/O/UV (i.e., infrared/optical/ultraviolet) “high-contrast direct imaging mission with a target off-axis inscribed diameter of approximately 6 meters” which resembles but does not coincide with the previous HabEx (Habitable Exoplanet Observatory, Gaudi et al., 2020) and LUVOIR (Large UV/Optical/IR Surveyor, The LUVOIR Team, 2019) concepts. In recent months, efforts have ramped up to perform architecture trade studies to eventually define an HWO concept, and the community has been heavily involved in proposing relevant science cases for the new Observatory. This is a critical task at this point in time since, as stated in the Decadal, “implementation should start in the latter part of the decade [2022-2032], with a target launch in the first half of the 2040s”, pending an independent review on the scientific and technological maturity. As the architecture for HWO is being developed, it is crucial to optimize the observing strategies for efficient target identification and characterization so that the mission requirements can be matched to the planned observations.
@@
A main exoplanet-related requirement for an HWO-like mission is to observe 30 - 50 terrestrial planets around nearby FGK stars over the span of 5 mission years, though the potential of a serviceable observatory is also discussed and would extend and enhance the nominal mission. Even with state-of-the-art coronagraphs, the spectroscopic characterization of planets whose flux is 10 billion times smaller than their host star (approximately the contrast we would need to detect an Earth-like planet around the Sun) would require days of observation (see e.g. Ulses et al., 2025). Furthermore, the presence of a coronagraph limits us to only observe a portion of the wavelength range at a time, thus increasing the required observing time for a full characterization of the planetary spectrum from the UV (i.e., ultraviolet) to the NIR (i.e. near-infrared).
@@
It is therefore essential to define optimal observing strategies for efficiently identifying which planetary discoveries are most likely to be habitable, and therefore of highest priority for spectroscopic follow-up. For the best targets, HWO will acquire spectroscopic data on the full wavelength range, and the current strategy baselined for target prioritization is through spectroscopy “decision tree” observation sequences (Young et al., 2024) which prioritize key molecules to detect that can distinguish different planet types. Bayesian retrievals have been used to identify what is the minimum and optimal wavelength ranges to perform observations for a variety of planetary atmospheres (see e.g., Feng et al., 2018; Damiano and Hu, 2022; Damiano et al., 2023; Latouf et al., 2023; Krissansen-Totton et al., 2025; Ulses et al., 2025); these observational requirements have recently been used when simulating exo-Earth survey yields (Morgan et al., 2024a, b; Stark et al., 2024a, b).
@@
The potential of colors and broadband photometry to qualitatively characterize exoplanets is not new. Traub (2003) argued that colors and low-resolution spectra in the VIS range would allow us to characterize atmospheric and surface spectral signatures, as well as determine physical parameters. On the other hand, Cahoy et al. (2010) performed a study on the albedo of gas giants and highlighted that phase-dependence and the presence of clouds could make this less straightforward and multiple observations would be needed to determine cloud distributions and orbital parameters. Krissansen-Totton et al. (2016) empirically calculated the optimal photometric bins to differentiate an Earth from a variety of other planetary classes. To achieve this, they identified the filters that maximized the distance in color-color space between the Earth and every other modeled planet. Batalha et al. (2018) explored the classification of gas giants through color information obtained with potential filter options for the Nancy Grace Roman Space Telescope, and the relation with metallicity, temperature, clouds, and phases. The authors performed a supervised classification multivariate analysis to assess if photometry would be sufficient to retrieve the underlying physics of an exoplanet. The analysis was done every combination of two up to five filters and found that at least three filters are needed to provide any classification. Smith et al. (2020) calculated a grid of potential rocky atmospheres and evaluated the potential for multiband photometry to effectively distinguish different classes of planetary atmospheres. They found the technique effective for atmospheres exhibiting strong absorption lines, but more challenging for Earth-like abundances and requiring large apertures to improve results.
Beyond the challenges of degeneracies in planetary atmospheres, distinguishing an Earth from other planetary types is made even more difficult when considering orbital geometry effects and unknown phase. As shown in Roberge and Seager (2018), the reflected light spectra of different planets span orders of magnitude when positioned at the same semi-major axis and phase. However, degeneracies can appear when altering these two parameters, and atmospheric features might become less distinguishable. Specifically, Roberge and Seager (2018) highlighted a scenario where a Neptune-like planet and an Earth would hardly be differentiable with each other at a given phase, especially at lower resolutions and within the average expected noise for current instrumentation. Given the limited observing time available for future direct imaging missions, developing strategies to efficiently distinguish between these planetary types becomes crucial for prioritizing truly habitable worlds over larger, uninhabitable planets that happen to fall within similar brightness ranges.
@@
We provide methodology and results for every step in the algorithm: Section III describes the orbital modeling; Section IV deals with the radiative transfer flux calculations; Section V provides details on the Exposure Time Calculator for HWO and the noise simulations; Section VI explores the best filter combinations for a successful discrimination between an Earth and Neptune-like planets. We discuss the implications and limitations of this study in Section VII, and we provide a summary in Section VIII.
II Methodology Overview
@@
We test this algorithm considering the Earth on a face-on orbit (quadrature phase) compared to two template Neptune scenarios (which we call “Warm” and “Cold”). We will expand on different phases and more realistic scenarios in an upcoming study (see Section VII.4).
III Orbital Calculations
@@ III.1 Results
In Figure 2 we show how orbits that satisfy condition 1) are distributed in semi-major axis space. We highlight in green the models that are within the commonly assumed boundaries for the habitable zone (HZ; Kopparapu et al., 2013), and in red are the ones outside the HZ (note that planets interior to the HZ would have smaller angular separations than our assumed 0.1 arcsec and thus would not satisfy the orbital condition). We notice that about 65% of the orbits that fit the assumed target projected separation are within the HZ. A non-negligible 35% of the compatible orbits lie well beyond the HZ, with most within 5 AU. If we assume no prior knowledge on the observed system, we would have no way of knowing from the first orbit which of these configurations is the correct one. While we could assume that it is more likely that a planet at the prescribed 0.1 arcsec separation is explained by an orbit within the HZ, scenarios including planets outside the HZ cannot be excluded.
It must be noted that these results are highly dependent on the uncertainty of the projected separation. For this study, we made an arbitrary choice to ensure that the potential Earth would be located within the HZ of its star (which explains the higher percentage of planets in this region), but in reality these requirements might be more relaxed, especially for more distant targets. We also assumed that a planet could be potentially found at any inclination or semi-major axis, disregarding any information that could come from planet population occurrence rates or precursor observations. We address the limitations of these assumptions in Section VII.4.
IV Flux Calculation
@@
Atmosphere Reference Kofman et al. (2024) Payne et al. (2025) Hu (2019)
Figure 3: Temperature and abundance profiles for template Earth and Neptunes atmospheres. Left: Earth atmosphere from Kofman et al. (2024) Center: Cold Neptune atmosphere from Payne et al. (2025) Right: Warm Neptune atmosphere from Hu (2019).
To calculate planetary reflectance spectra, we use the Planetary Spectrum Generator (PSG) [^1] https://psg.gsfc.nasa.gov (Villanueva et al., 2018). PSG is a radiative transfer model suite hosting tools that are widely used to model not only confirmed exoplanets, but also Solar System planets and cometary objects. PSG can compute 1D and 3D radiative transfer models from any configuration and allows the flexibility to choose instrumental parameters for nearly any observatory. It can ingest billions of spectral lines of almost 1000 species across its spectral repositories (HITRAN, JPL, CDMS, GSFC-Fluor, ExoMol), allowing the user to produce detailed high resolution spectra.
@@
We therefore enabled the “disk sub-sampling” mode in PSG to produce planetary reflectances at different phase angles. In this mode, PSG divides the disk of the planet into multiple concentric regions of common incidence and emission angles, each with a specific weight based on the disk percentage the region occupies. For every bin, the radiative transfer equation is solved and the output flux is a weighted mean of the fluxes of each region. In our study, we assumed a sampling of $N=5$, which creates five bins for the incidence angle and five more for the emission angle. As these two angles change with phase, the total amount of regions that are created (as the possible combinations of the incidence and emission angle) varies, up to around 25 ($\approx N^{2}$). Saxena et al. (2021) discusses PSG’s disk sub-sampling algorithm in more depth.
@@
The algorithm we developed for this study is interfaced with an instance of PSG installed locally via a Docker container [^2] https://docs.docker.com/get-started/docker-concepts/the-basics/what-is-a-container/, enabling repeated and rapid calls to PSG (see Villanueva et al., 2018, 2022, for more information on how to implement a “dockerized” PSG version). For every model to be calculated, the algorithm produces a configuration file that describes the desired input (including the atmospheric profile, the desired opacities to be used, the surface and aerosol scattering parameters, the stellar class and the orbital parameters) as well as the format of the desired output, and then reads in the results of the calculation.
@@
For the reference scenario (Earth at quadrature), we calculate the flux of an Earth twin orbiting a Sun-twin star at 1 AU. For the atmospheric profile for a standard Earth-like planet, we use an atmosphere based on Kofman et al. (2024), modeled using the Modern-Era Retrospective analysis for Research and Applications (MERRA-2) 3D climate database [^3] https://gmao.gsfc.nasa.gov/gmao-products/merra-2/. The 3D distribution of temperature, pressure, clouds and relevant molecules from MERRA-2 were ingested into the Global Emission Spectrum (GlobES, Fauchez et al., 2025) module of the Planet Spectrum Generator (PSG Villanueva et al., 2018)). In the MERRA-2 reanalysis database, the 3D abundance and presence of clouds was parameterized in a 2D cloud fraction field. PSG evaluates the spectra from the MERRA-2 input by running two spectral calculations, one considering a clear sky and one assuming a fully clouded atmosphere according to the ingested cloud particle abundances. The two spectra are combined following the cloud fraction.
@@
Surface coverage was adopted from the Moderate Resolution Imaging Spectroradiometer (MODIS) [^4] https://modis.gsfc.nasa.gov using five different surface reflectances from the United States Geological Survey (USGS) [^5] https://www.usgs.gov/labs/spectroscopy-lab/usgs-spectral-library. The data from this source were validated with narrow band photometric observations from DSCOVR (between 0.318-0.780 microns), as well as several other observations of Earth (Payne et al., 2025). We show in Figure 3 and Table 1 the atmospheric and temperature profiles for the Earth, as well as some relevant input parameters.
@@
We then calculate the flux of all compatible orbit configurations from Section III, assuming two different Neptune scenarios: if the semi-major axis is less than 3.5 AU (1167 out of 1435 Neptunes), then we assume a warm Neptune atmosphere; otherwise (268 out of 1435 Neptunes), we assume a cold Neptune atmosphere. The choice of the 3.5 AU threshold is motivated by estimates for the snow line (Martin and Livio, 2012, see e.g.,), which is a sensible physical separation between classes of Neptune-like planets with different atmospheric compositions and thermal structures. Beyond this boundary, a water or ammonia cloud deck is expected above $\sim$1 bar (Hu, 2019; Cahoy et al., 2010).
@@
We use fixed atmospheric templates for each Neptune type to isolate the effects of orbital geometry (phase angle and distance) on observed fluxes, allowing us to focus on how orbital parameters alone can create spectral confusion. We discuss the implications of this assumption in Section VII.4 and we will consider self-consistent chemistry in a future work.
@@
As warm Neptune template, we use one of the models calculated in Hu (2019). We selected a model for an H2/He-dominated warm Neptune at a semi-major axis 1.7 AU (a good medium between the edges of the warm Neptune subgrid), internal heat flux $T_{int}=60\ K$ to mimic a Neptune-like planet, gravity of $10\ m/s^{2}$ (close to that of Neptune), and 10x solar metallicity (similar to the metallicity found in observed warm Neptune-sized planets). The models were computed using ExoREL, which calculates pressure-temperature profiles using a grey-atmosphere approximation and taking into account the condensation of water and ammonia clouds in a self-consistent way.
The cold Neptune template was derived from all available measurements of the Solar System’s Neptune (Payne et al., 2025). The cold Neptune continuum is based on data from the Hubble Space Telescope (Courtin, 1999), the European Southern Observatory (Karkoschka, 1998) and the NASA Infrared Telescope Facility (Irwin et al., 2022). The collection of observations provide the composite reflectance spectrum that was used as a reference to develop the best-fit reflectance model in PSG by adjusting atmospheric absorbers, aerosols, and cloud features to match the reference spectrum; additional details are available in Payne et al. (2025).
We simulate the “discovery” flux measurement by calculating the average photometric flux of the Earth at 500 nm assuming 20% bandwidth $F_{500,20\%}$. This translates into averaging the flux in the interval $500\cdot(1-10\%),500\cdot(1+10\%)]$ nm =[450,550]=[450,550] nm. Since the currently-assumed signal-to-noise ratio ($S/N$) for detection is 7 (Nemati et al., 2020) at the detection wavelength range, the errorbar associated with the measurement $\Delta F_{500,20\%}=F_{500,20\%}/7$. The flux of a Neptune is then considered compatible if the averaged Neptune flux over the discovery bandpass lies within $F_{500,20\%}\pm\Delta F_{500,20\%}$.
IV.1 Results